a railway half ticket costs half the full fare and the reservation charge for the distance covered. for a journey of 12km the charge paid is 89 and for a journey of 20km the charge paid is 145. what will a person have to pay for travelling a distance of 30km?
Answer by rwm(914)

h=t/2+r*d
89=t/2+r*12
145=t/2+r*20
x=t/2+r*30
r=7
t=10
x=215
30 km costs 215